OpenStudy (acal21):

((-1)^n*n)/(3+2n^2), this series converges, but i cant quite find the limit. need to find the limit

OpenStudy (dumbcow):

pretty sure limit is 0 (-1)^n simply alternates between 1 and -1 \[\lim_{n \rightarrow \infty} \frac{n}{3+2n^2} = 0\]

ganeshie8 (ganeshie8):

are you really trying to find the limit of series ? @acal21

ganeshie8 (ganeshie8):

By alternating series test, the series converges. I'm not sure if you really want to find the sum

OpenStudy (acal21):

YES CAUSE THIS SERIES CONVERGES , THE PROBLEMS ASK ME TO FIND THE LIMIT IF IT CONVERGES